OpenStudy (anonymous):

FAN & MEDAL! Can anyone help me with ratios please? I have a few questions for an online exam that are part of my end of the year grade. They are 6th grade. Please help! I am horrible with ratios, I cant take this test again!

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Alice and Brian have cookies in a ratio of 4:3. Brian and Cori have cookies in a ratio of 6:4. What is the ratio of Alice’s cookies to Cori’s cookies? A. 8 to 4 B. 4 to 8 C. 3 to 6 D. 6 to 4

OpenStudy (isaidavila):

I would say is 8 to 4

OpenStudy (anonymous):

definity not c or d

OpenStudy (anonymous):

thanks, I need proof that its the right one though.

OpenStudy (anonymous):

So either A. or B.

OpenStudy (anonymous):

What do you think?

OpenStudy (isaidavila):

A/B=4/3 B/C=6/4 But look at B in the fir ratio , he has 3 but in the second ve has 6 so multiply numerator and denominator( of the first) by 2 so we get A/B=8/6 B/C=6/4 So A/C=8/4

OpenStudy (anonymous):

Thank you soooooo much!

OpenStudy (anonymous):

could I ask you a few more?

OpenStudy (isaidavila):

Yeah

OpenStudy (anonymous):

Let me pull it up

OpenStudy (anonymous):

This table shows the numbers of books, by type, checked out from the school library on Monday. Book Checkout Book Type Number of Books mystery 24 nonfiction 18 adventure 12 humor 16 What is the ratio of mystery books checked out to nonfiction books checked out? A. 4 to 3 B. 1 to 2 C. 3 to 4 D. 2 to 1

OpenStudy (isaidavila):

Just divide the corresponding numbers : 24/18 = 4/3

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

so A.?